If area of triangle is 2√2 unit.then least value of s (semi perimeter) is.........

Answer by richard1234(7193)

Intuitively, the perimeter should be maximized when the triangle is equilateral, but it seems much harder to prove. By AM-GM, we have



The RHS is a constant, and the LHS is equal to s/2 (since a+b+c = 2s). However the equality case cannot occur since this would imply s = s-a = s-b = s-c, which gives a = b = c = 0.